Could you please help me solve for X and state the restrictions on the variables: bx+a=dx+c ? Every time I tried I always ended up with my x's being canceled because I couldn't figure out what I was doing wrong when I divided by both b and d.
Answer by richwmiller(17219)

get all x's on one side and factor
bx+a=dx+c
bx-dx=c-a
x(b-d)=(c-a)
x=(c-a)/b-d)
b can not be equal to d
so that b-d can't be zero